OpenStudy (anonymous):

(y cubed) -1/6 power (simplify)

OpenStudy (anonymous):

1/y^(1/2)

OpenStudy (anonymous):

\[(y^3)^{-1/6}=\frac{1}{y^{3*\frac{1}{6}}}=\frac{1}{y^{1/2}}\]or in simple radical form\[=\frac{\sqrt{y}}{y}\]